# Non-Markovian Discrete Diffusion with Causal Language Models

### v2

@@ Abstract
Discrete diffusion models offer a flexible, controllable approach to structured sequence generation, yet they still lag behind causal language models in expressive power. A key limitation lies in their reliance on the Markovian assumption, which restricts each step to condition only on the current state, leading to potential uncorrectable error accumulation. In this paper, we introduce CaDDi, a discrete diffusion model that conditions on the entire generative trajectory, thereby lifting the Markov constraint and allowing the model to revisit and improve past states. By unifying sequential (causal) and temporal (diffusion) reasoning in a single non-Markovian transformer, CaDDi also treats standard causal language models as a special case and permits the direct reuse of pretrained LLM weights with no architectural changes. Empirically, CaDDi outperforms state-of-the-art discrete diffusion baselines on natural-language benchmarks, substantially narrowing the remaining gap to large autoregressive transformers.
@@ 2 General Discrete Diffusion
where $\operatorname{Cat}(\cdot;\,\pi)$ denotes the categorical distribution with parameter $\pi$, and $[\mathbf{Q}_{t}]_{ij}$ gives the probability of transitioning from state $i$ to state $j$ at time $t$. Due to the Markovian assumption, the marginal distribution at any timestep can be computed in closed form:
@@
Figure 1: (a). Inference paradigm for a standard causal language model versus CaDDi-AR. In CaDDi-AR, each timestep first autoregressively denoises the tokens into $\widetilde{\mathbf{x}}_{0}$, then re-applies noise via the diffusion kernel to obtain A traditional autoregressive model emerges as the special case of $T=1$, which can be adapted to discrete diffusion by fine-tuning. (b). Extending 1D to 2D Rotary Positional Encoding. Standard rotary encodings for token positions are seamlessly generalized to also encode diffusion timesteps, remaining fully backward-compatible with existing language model architectures.
@@ 3.2 Non-Markovian Forward Process
[…]
4:Predict $\widetilde{\mathbf{x}}_{0}\sim\mu_{\theta}\left(\mathbf{x}_{0}\mid\mathbf{x}_{t:T},t\right)$
5:Sample $\mathbf{x}_{t-1}\sim q\left(\mathbf{x}_{t-1}\mid\widetilde{\mathbf{x}}_{0}\right)$EndFor
[…]
@@ 3.3 Non-Markovian Inference Process
Notice that it lift the constraint introduced by $\mathbf{x}_{t:T}$ in the posterior form and delegating all the flexibility for the denoiser model $\mu_{\theta}$. As illustrated in Fig. 2 (right), inference proceeds in an autoregressive manner: at each timestep, inference proceeds in an autoregressive manner: at each timestep $t$, we first predict a clean estimate $\widetilde{\mathbf{x}}_{0}\sim\mu_{\theta}\left(\mathbf{x}_{t:T},t\right)$, then sample $\mathbf{x}_{t-1}$ using the forward corruption kernel $q\left(\mathbf{x}_{t-1}\mid\widetilde{\mathbf{x}}_{0}\right)$. This loop continues backward through time until the final data sample $\mathbf{x}_{0}$ is recovered.
@@ 3.4 Evidence Lower Bound
where $\mathcal{L}_{T}=\sum_{t=2}^{T}\mathbb{E}_{q(\mathbf{x}_{t:T}\mid\mathbf{x}_{0})}\operatorname{KL}(q(\mathbf{x}_{t-1}\mid\mathbf{x}_{0})\|p_{\theta}\bigl{(}\mathbf{x}_{t-1}\mid\mathbf{x}_{t:T})$. The second term is constant for most forward kernels. For specific kernel choices (e.g. absorbing or uniform), the ELBO simplifies further:
@@ 4.1 Unified Sequential and Temporal Modeling
While it’s straightforward to use a decoder-only causal model for modeling the sequential and temporal dependncy of latent chains. One of the issue is that modern causal language model typically encodes only the sequential dimension, via rotary positional encodings [53]. However, for non-Markovian discrete diffusion, we must capture not just the standard token-level sequence but also a temporal dimension corresponding to diffusion timesteps. To address this, we extend the original 1D rotary scheme to a 2D variant.
@@ 4.2 CaDDi-AR: Factorization over Token Space
As dicussed in [23, 63, 62], independent factorization over dimensions in one reverse step can make $p_{\theta}\left(\mathbf{x}_{t-1}\mid\mathbf{x}_{t:T}\right)$ neglecting the token dependence and fail to match the true posterior $q\left(\mathbf{x}_{t-1}\mid\mathbf{x}_{t:T}\right)$. A straightforward solution following [23] is to further factorizing over token space - leading to a token-level autoregression $p_{\theta}\left(\mathbf{x}_{t-1}\mid\mathbf{x}_{t:T}\right)=\prod_{i=0}^{L}p_{\theta}\left(\mathbf{x}_{t-1}^{i}\mid\mathbf{x}_{t-1}^{0:i-1},\mathbf{x}_{t:T}\right)$ and next-token prediction loss. We denote this varient as CaDDi-AR. Notice it naturally fits our decoder-only causal language model, where the sampling process essentially correpond to autoregressive generation with historical trajectory $\mathbf{x}_{t:T}$ serving as "prompt". The token-level autoregression decomposing enables CaDDi-AR a more fine-grained granulairity in per-step generation.
@@ 5 Experiments
We evaluate CaDDi’s generative capabilities on the One Billion Words dataset [7], a large-scale natural language corpus comprising over 30 million English sentences of varying lengths. We follow the tokenization and training setup introduced in DiffusionBERT [26]. For UDLM, we use the pretrained weights provided by the authors. For other baselines, we retrain the models using their official codebases, as checkpoints are not publicly available.
@@
Following prior work [2, 48], we trained a discrete diffusion model on short text chunks of length 256 from the text8 dataset. We use the same dataset split as in previous studies, training on the training set and reporting performance on the test set using the standard bits-per-character (BPC) metric. The BPC is defined as: $-\frac{1}{L}\sum_{i=1}^{L}\log_{2}p\left(\mathbf{x}_{i}\right)$. Note that the bound of log-likelihood depends on the discretization of the diffusion schedule. To enable fair comparisons, we report the number of discretization steps used for each diffusion-based model.
We omit BPC results for CaDDi-AR due to the additional approximation error of likelihood introduced by its token-level autoregressive decomposition. As shown in Table 2, under the same discretization setting (64 steps), CaDDi outperforms other discrete diffusion models. While any-order autoregressive models can be viewed as a special case of discrete diffusion with 256 steps, our 64-step model already approaches their performance.
@@ 30. Guidelines:
  33. Answer: [Yes]
  34. Justification: Experiment results with error bar is provided if applicable.
@@ Appendix A Proof
[…]
@@ A.1.1 Forward–process definitions
[…]
@@ A.1.2 Mutual–information computations and decay measure
    Un-masked: with probability $1-\alpha_{t}^{*}$ the token is uncorrupted and $\mathbf{x}_{t}=\mathbf{x}_{0}$.
@@
Only two cases have non-zero probability:
@@
    If $Z=0$ (at least one un-masked coordinate, probability $1-\prod\alpha_{\tau}$), every un-masked coordinate equals $\mathbf{x}_{0}$; thus $\mathbf{x}_{0}$ is known exactly and $H(\mathbf{x}_{0}\mid\mathbf{x}_{t:T})=0$.
@@
Most existing studies on absorbing-state Markovian diffusions [2, 37, 44] adopt the linear cumulative–masking curve $\alpha_{t}^{\mathrm{lin},*}=t/T,\;t=0,\dots,T.$ For a fair comparison by default we employ a non-Markovian independent schedule ${\alpha}^{\mathrm{lin}}$ that reproduces exactly the same marginals in our experiments, i.e. satisfies $\alpha_{t}^{\mathrm{lin},*}=\prod_{\tau=t}^{T}\alpha_{\tau}^{\mathrm{lin}}.$ Using the backward recursion $\alpha_{t}=\alpha_{t}^{*}/\alpha_{t+1}^{*}$ (Sec. A.1) gives the closed-form
@@ A.2 Derivation of the Non-Markovian Evidence Lower Bound (ELBO)
[…]
@@ C.1.2 Trajectory Re-composition
$\mathbf{x}_{t}^{\curvearrowright}$ $:=\mathbf{x}_{t}\oplus\mathbf{x}_{t+1}\oplus\cdots\oplus\mathbf{x}_{T}$ (30)
$\Gamma_{\mathrm{rec}}\left(\mathbf{x}_{1:T}\right)$ $:=\operatorname{FlattenTime}\left(\mathbf{x}_{t}^{\curvearrowright},\mathbf{x}_{t+1}^{\curvearrowright},\ldots,\mathbf{x}_{t+m-1}^{\curvearrowright}\right)$

### v3

@@ Abstract
Discrete diffusion models offer a flexible, controllable approach to structured sequence generation, yet they still lag behind causal language models in expressive power. A key limitation lies in their reliance on the Markovian assumption, which restricts each step to condition only on the current state, leading to potential uncorrectable error accumulation. In this paper, we introduce CaDDi (Causal Discrete Diffusion Model), a discrete diffusion model that conditions on the entire generative trajectory, thereby lifting the Markov constraint and allowing the model to revisit and improve past states. By unifying sequential (causal) and temporal (diffusion) reasoning in a single non-Markovian transformer, CaDDi also treats standard causal language models as a special case and permits the direct reuse of pretrained LLM weights with no architectural changes. Empirically, CaDDi outperforms state-of-the-art discrete diffusion baselines on natural-language benchmarks, substantially narrowing the remaining gap to large autoregressive transformers.
@@ 2 General Discrete Diffusion
where $\operatorname{Cat}(\cdot;\,\pi)$ denotes the categorical distribution with parameter $\pi$, and $\mathbf{Q}_{t}]_{ij}$ gives the probability of transitioning from state $i$ to state $j$ at time $t$. Due to the Markovian assumption, the marginal distribution at any timestep can be computed in closed form:
@@
Figure 1: (a). Inference paradigm for a standard causal language model versus CaDDi-AR. In CaDDi-AR, each timestep first autoregressively denoises the tokens into $\widetilde{\mathbf{x}}_{0}$, then re-applies noise via the diffusion kernel to obtain ${\mathbf{x}}_{t-1}$. A traditional autoregressive model emerges as the special case of $T=1$, which can be adapted to discrete diffusion by fine-tuning. (b). Extending 1D to 2D Rotary Positional Encoding. Standard rotary encodings for token positions are seamlessly generalized to also encode diffusion timesteps, remaining fully backward-compatible with existing language model architectures.
@@ 3.2 Non-Markovian Forward Process
3: Initialize $\mathbf{x}_{T}\sim q(\mathbf{x}_{T})$ as the noisy input at the final timestep
4: for $t=T$ down to 11 do
5:   Predict $\widetilde{\mathbf{x}}_{0}\sim\mu_{\theta}\left(\mathbf{x}_{0}\mid\mathbf{x}_{t:T},t\right)$
6:   Sample $\mathbf{x}_{t-1}\sim q\left(\mathbf{x}_{t-1}\mid\widetilde{\mathbf{x}}_{0}\right)$
7: return $\mathbf{x}_{0}$
@@ 3.3 Non-Markovian Inference Process
Notice that it lift the constraint introduced by $\mathbf{x}_{t:T}$ in the posterior form and delegating all the flexibility for the denoiser model $\mu_{\theta}$. As illustrated in Fig. 2 (right), inference proceeds in an autoregressive manner: at each timestep, we first predict a clean estimate $\widetilde{\mathbf{x}}_{0}\sim\mu_{\theta}\left(\mathbf{x}_{t:T},t\right)$, then sample $\mathbf{x}_{t-1}$ using the forward corruption kernel $q\left(\mathbf{x}_{t-1}\mid\widetilde{\mathbf{x}}_{0}\right)$. This loop continues backward through time until the final data sample $\mathbf{x}_{0}$ is recovered.
@@ 3.4 Evidence Lower Bound
where $\mathcal{L}_{T}=\sum_{t=2}^{T}\mathbb{E}_{q(\mathbf{x}_{t:T}\mid\mathbf{x}_{0})}\operatorname{KL}(q(\mathbf{x}_{t-1}\mid\mathbf{x}_{0})\|p_{\theta}\bigl(\mathbf{x}_{t-1}\mid\mathbf{x}_{t:T}))$. The second term is constant for most forward kernels. For specific kernel choices (e.g. absorbing or uniform), the ELBO simplifies further:
@@ 4.1 Unified Sequential and Temporal Modeling
While it’s straightforward to use a decoder-only causal model for modeling the sequential and temporal dependency of latent chains. One of the issue is that modern causal language model typically encodes only the sequential dimension, via rotary positional encodings [53]. However, for non-Markovian discrete diffusion, we must capture not just the standard token-level sequence but also a temporal dimension corresponding to diffusion timesteps. To address this, we extend the original 1D rotary scheme to a 2D variant.
@@ 4.2 CaDDi-AR: Factorization over Token Space
As dicussed in [23, 63, 62], independent factorization over dimensions in one reverse step can make $p_{\theta}\left(\mathbf{x}_{t-1}\mid\mathbf{x}_{t:T}\right)$ neglecting the token dependence and fail to match the true posterior $q\left(\mathbf{x}_{t-1}\mid\mathbf{x}_{t:T}\right)$. A straightforward solution following [23] is to further factorizing over token space - leading to a token-level autoregression $p_{\theta}\left(\mathbf{x}_{t-1}\mid\mathbf{x}_{t:T}\right)=\prod_{i=0}^{L}p_{\theta}\left(\mathbf{x}_{t-1}^{i}\mid\mathbf{x}_{t-1}^{0:i-1},\mathbf{x}_{t:T}\right)$ and next-token prediction loss. We denote this varient as CaDDi-AR. Notice it naturally fits our decoder-only causal language model, where the sampling process essentially correpond to autoregressive generation with historical trajectory $\mathbf{x}_{t:T}$ serving as "prompt". The token-level autoregression decomposing enables CaDDi-AR a more fine-grained granularity in per-step generation.
@@ 5 Experiments
We evaluate CaDDi’s generative capabilities on the One Billion Words dataset (LM1B) [7], a large-scale natural language corpus comprising over 30 million English sentences of varying lengths. We follow the tokenization and training setup introduced in DiffusionBERT [26]. For UDLM, we use the pretrained weights provided by the authors. For other baselines, we retrain the models using their official codebases, as checkpoints are not publicly available.
@@
Following prior work [2, 48], we trained a discrete diffusion model on short text chunks of length 256 from the text8 dataset. We use the same dataset split as in previous studies, training on the training set and reporting performance on the test set using the standard bits-per-dimension (BPD) metric. The BPD is defined as: $\frac{1}{L}\sum_{i=1}^{L}\log_{2}p\left(\mathbf{x}_{i}\right)$. Note that the bound of log-likelihood depends on the discretization of the diffusion schedule. To enable fair comparisons, we report the number of discretization steps used for each diffusion-based model.
We omit BPD results for CaDDi-AR due to the additional approximation error of likelihood introduced by its token-level autoregressive decomposition. As shown in Table 2, under the same discretization setting (64 steps), CaDDi outperforms other discrete diffusion models. While any-order autoregressive models can be viewed as a special case of discrete diffusion with 256 steps, our 64-step model already approaches their performance.
@@ 6 Conclusion
Acknowledgement
The authors thank collaborators and contributors from across institutions for their invaluable support and insights throughout this project. This work was supported in part by the National Institutes of Health (NIH) grant R35GM143072–01, the National Science Foundation (NSF) grant IIS Division of Information & Intelligent Systems 2443528, and the Yale Colton Center Award, all awarded to Prof. David van Dijk; and by the National Science Foundation (NSF) grants IIS Division of Information & Intelligent Systems 2403317 and CNS Division of Computer and Network Systems 2431504, as well as the Envisioning Artificial Intelligence at Yale 2025 initiative from the Yale Office of the Provost, awarded to Prof. Rex Ying.
@@ 30. Guidelines:
  33. Answer: [No]
  34. Justification: Error bars are not reported because it would be too computationally expensive
@@ Appendix A Proof
Throughout this appendix we work with a single–token random variable that takes values in the augmented vocabulary $\mathcal{V}_{+}:=\mathcal{V}\cup\{\texttt{[MASK]}\}$. The distinguished absorbing symbol is denoted by $m\in\mathcal{V}_{+}$ and we write $e_{m}$ for the corresponding standard basis vector.
@@ A.1.1 Forward–process definitions
$\mathbf{x}_{s}\mid\mathbf{x}_{0}\overset{\text{i.i.d.}}{\sim}(1-\alpha_{s})\delta_{\mathbf{x}_{0}}+\alpha_{s}\delta_{m},\qquad s=1,\dots,T.$ (12)
@@ A.1.2 Mutual–information computations and decay measure
    Un‑masked: with probability $1-\alpha_{t}^{*}$ the token is uncorrupted and $\mathbf{x}_{t}=\mathbf{x}_{0}$.
@@
Only two cases have non‑zero probability:
@@
    If $Z=0$ (at least one un‑masked coordinate, probability $1-\prod\alpha_{\tau}$), every un‑masked coordinate equals $\mathbf{x}_{0}$; thus $\mathbf{x}_{0}$ is known exactly and $H(\mathbf{x}_{0}\mid\mathbf{x}_{t:T})=0$.
@@
Most existing studies on absorbing‑state Markovian diffusions [2, 37, 44] adopt the linear cumulative–masking curve $\alpha_{t}^{\mathrm{lin},*}=t/T,\;t=0,\dots,T.$ For a fair comparison by default we employ a non‑Markovian independent schedule ${\alpha}^{\mathrm{lin}}$ that reproduces exactly the same marginals in our experiments, i.e. satisfies $\alpha_{t}^{\mathrm{lin},*}=\prod_{\tau=t}^{T}\alpha_{\tau}^{\mathrm{lin}}.$ Using the backward recursion $\alpha_{t}=\alpha_{t}^{*}/\alpha_{t+1}^{*}$ (Sec. A.1) gives the closed‑form
@@ A.2 Derivation of the Non-Markovian Evidence Lower Bound (ELBO)
$\log p_{\theta}\left(\mathbf{x}_{0}\right)=\log\int p_{\theta}\left(\mathbf{x}_{0},\mathbf{x}_{1:T}\right)d\mathbf{x}_{1:T}=\log\int\frac{p_{\theta}\left(\mathbf{x}_{0},\mathbf{x}_{1:T}\right)}{q\left(\mathbf{x}_{1:T}\mid\mathbf{x}_{0}\right)}q\left(\mathbf{x}_{1:T}\mid\mathbf{x}_{0}\right)d\mathbf{x}_{1:T}$
@@ C.1.2 Trajectory Re-composition
$\mathbf{x}_{t}^{\curvearrowright}$ $=\mathbf{x}_{t}\oplus\mathbf{x}_{t+1}\oplus\cdots\oplus\mathbf{x}_{T}$ (30)
$\Gamma_{\mathrm{rec}}\left(\mathbf{x}_{1:T}\right)$ $=\operatorname{FlattenTime}\left(\mathbf{x}_{t}^{\curvearrowright},\mathbf{x}_{t+1}^{\curvearrowright},\ldots,\mathbf{x}_{t+m-1}^{\curvearrowright}\right)$